Find the heat transferred from the nitrogen to the surroundings. Use cv for nitrogen = 0.75 kJ/kg-K.


1
Expert's answer
2021-05-04T06:47:58-0400

Full question must be like this :-


Nitrogen gas having a mass of 0.8 kg at 110 kPa and 50℃ is contained in a cylinder. The piston is moved compressing nitrogen until the pressure becomes 1.1 MPa and temperature becomes 200℃. The work done during the process is 50 kJ. Find the heat transferred from the nitrogen to the surroundings. Use 𝑐𝑣 for nitrogen = 0.75 kJ/kg-K.


Answer :-


"\nu=\frac mM=\frac{0.8}{0.028}=28.6~mole"


"\gamma = 1 + \frac{R}{C_v}=1 + \frac{8.31}{750}= 1.011"


"A= {vrT_1\over \gamma -1}(1-(\frac{p_2}{p_1}.\frac{T_2}{T_1})^{\gamma -1}) = 150 kJ ,"



"Q=A-A'=100~kJ."